Required information

[The following information applies to the questions displayed below.]

Riverview Company's budget for the coming year includes $6,000,000 for manufacturing overhead, 50,000 hours of direct labor, and 250,000 hours of machine time.

If Riverview applies overhead using a predetermined rate based on machine-hours, what amount of overhead will be assigned to a unit of output which requires 0.5 machine hours and 0.25 labor hours to complete?

Multiple Choice

  • $12.00

  • $16.00

  • $ 6.00

  • $20.00
Answer #1

Predetermined Overhead rate

= Estimated overhead/Estimated machine hours

= 6,000,000/250,000 = 24

Overhead assigned to a unit of output

= 24*0.5

= 12

Option A is the answer